# Dark tranquility: Overdamped quasibound states inside a Schwarzschild black hole

### v1

@@ V Discussion
The interior of a Schwarzschild BH is traditionally regarded as a region where all physical fields are inexorably driven into the singularity. In this work, we have demonstrated that axial gravitational perturbations in Schwarzschild BH interiors can give rise to OQBSs that linger without ever reaching the central singularity. Specifically, using Kruskal–Szekeres coordinates, we identified and analyzed these modes, showing that their Regge–Wheeler master wavefunctions remain finite at the future horizon while decaying before the singularity. This result runs counter to the common assumption that all disturbances inside the BH are ineluctably drawn into the singularity, pointing instead to a richer and more nuanced landscape of the dynamics within BH interiors.
@@
In addition to studying a spectral problem in the interior of a Schwarzschild BH, we have found that axial metric disturbances can hover between the horizon and the singularity via OQBSs, that retain some regularity. The results of this work apply to axial gravitational perturbations, only. For scalar or polar gravitational perturbations, the potential approaches -$\infty$ like $1/4r_{*}^{2}$ at the central singularity, which implies that the wavefunction vanishes like $\psi\sim(-r_{*})^{1/2}$ there. The vanishing of the wavefunction seems qualitatively similar to the case of axial perturbations, but the unbounded potential that diverges to -$\infty$, together with the lack of a potential well, presents a rather different problem from the axial case. From a numerical point of view, the polar case seems unstable against the formation of short wavelength oscillations in the large negative potential.
@@
Even though our analysis lies purely at the level of classical perturbation theory, there is a plurality of applications regarding quantum effects, semi-classical and eventually quantum gravity in BH interiors [157, 158]. Here we have obtained interior purely imaginary resonances. If appropriate regularity conditions are satisfied by the resonance wavefunctions, those may reflect decay rates or spatial modes relevant to infalling observers, semi-classical stress tensors, or even effective field theory on curved spacetime. Of course, these resonances do not couple in any physical way to the exterior and resulting observables but they might influence internal wave dynamics and late-time tails [159], how semi-classical fields diverge or remain finite near the singularity and possible holographic duals or quantum-gravity boundary conditions [160]. Thus, these resonances are meaningful in BH interior physics, especially if we consider how fields behave near the singularity or how quantum corrections might modify it. One might speculate that they could be useful for approaches like firewalls [161], and regular BHs [162, 163, 164] that lack a central singularity.
@@
In quantum gravity, especially in path integral or Euclidean approaches, boundary conditions at singularities or at finite radii are required [174]. Interior resonances could offer a natural selection mechanism for such boundary conditions, e.g., only allowing fields that are smooth or decay in a certain way at $r=0$ [175, 176]. These can even be viewed as “inner quantization conditions” [177], reminiscent of the Bohr-Sommerfeld conditions in semi-classical quantum mechanics [178]. In speculative models where BHs have dual descriptions [179, 180], e.g., fuzzballs [127, 181, 182], firewalls [161], and holographic duals [183, 184, 185], the interior degrees of freedom should encode information about microstates [186]. Interior resonant modes might act like a spectral fingerprint of interior microstructure, even if the actual microstates are unknown. So, even though these modes are not directly observable from outside the BH, they can influence the behavior of semi-classical fields, provide natural regularization or extension criteria, guide the construction of effective theories with modified interiors, or be incorporated into quantum corrections or path integrals, e.g., via mode sum regularization [187].
@@ Acknowledgements.
We thank Valentin Boyanov, Vitor Cardoso, Nicola Franchini, David Hilditch, Kostas D. Kokkotas, Adrien Kuntz, Jorge Patiño, Hannes Rüter and Sebastian Völkel for helpful discussions. This work was supported by the Israel Science Foundation, grant 531/22. K.D. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. R.B. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., through the ERC-Portugal program Project “GravNewFields”. This project has received funding from the European Union’s Horizon MSCA-2022 research and innovation programme “Einstein Waves” under grant agreement No. 101131233.
@@ B.2 Behavior under the Chandrasekhar-Darboux map
Thus, the fate of isospectrality can be decided solely from the requirement of regularity in the vicinity of $r=0$. If $\psi_{\rm RW}\sim r^{3}$, then
@@
Therefore, although the Regge-Wheeler and Zerilli equations are Darboux-related, their boundary-value problems are not isospectral in the BH interior. Hence, axial and polar modes inside Schwarzschild BHs do not share the same spectrum, as expected.

### v2

@@ V Discussion
The interior of a Schwarzschild BH is traditionally regarded as a region where all physical fields are inexorably driven into the singularity. In this work, we have demonstrated that axial gravitational perturbations in Schwarzschild BH interiors can give rise to OQBSs that linger temporarily, before decaying, as they drift towards the central singularity. Specifically, using Kruskal–Szekeres coordinates, we identified and analyzed these modes, showing that their Regge–Wheeler master wavefunctions remain finite at the future event horizon while decaying before the singularity. This result runs counter to the common assumption that all disturbances inside the BH are ineluctably drawn directly into the singularity, pointing instead to a richer and more nuanced landscape of the dynamics within BH interiors.
@@
In addition to studying a spectral problem in the interior of a Schwarzschild BH, we have found that axial metric perturbations can hover transiently between the horizon and the singularity via these OQBSs, that retain their regularity up to $r=0$. The results of this work apply to axial gravitational perturbations, only. For scalar or polar gravitational perturbations, the potential approaches -$\infty$ like $1/4r_{*}^{2}$ at the central singularity, which implies that the wavefunction vanishes like $\psi\sim(-r_{*})^{1/2}$ there. The vanishing of the wavefunction seems qualitatively similar to the case of axial perturbations, but the unbounded potential that diverges to -$\infty$, together with the lack of a potential well, presents a rather different problem from the axial case. From a numerical point of view, the polar case seems unstable against the formation of short wavelength oscillations in the large negative potential.
Figure 4: An OQBS as seen by geodesic observers. The red and green curves are the values from Fig. 3 along the red and green world lines, as functions of each observer’s proper time. The two curves are equal at the black circle, which corresponds to the white circle in Fig. 3.
An important reminder here is the fact that all physically-meaningful conclusions have been drawn from the wave dynamics of axial gravitational perturbations in the BH interior analyzed in Kruskal-Szekeres, instead of standard Schwarzschild coordinates. That being said, the OQBSs might be considered as “bound states in time” in a traditional Schwarzschild-coordinate perspective, since in the BH interior the conventional coordinate $r$, and thus $r_{*}$, is timelike. It is rather interesting, though, that the concept of such bound states has already been considered in other physical settings. For instance, Ref. [161] presents an analogous phenomenon in the context of Maxwell’s equations in a time-varying, spatially homogeneous medium. The equations used in [161] share a very similar structure with those studied here, with the respective effective potential arising from a time-dependent refractive index. This parallel reinforces the idea that such solutions can have a variety of meaningful physical interpretations.
@@
Even though our analysis lies purely at the level of classical perturbation theory, there is a plurality of applications regarding quantum effects, semi-classical and eventually quantum gravity in BH interiors [162, 163]. Here, we have obtained interior purely imaginary resonances. If appropriate regularity conditions are satisfied by the resonance wavefunctions, those may reflect decay rates or spatial modes relevant to infalling observers, semi-classical stress tensors, or even effective field theory on curved spacetime. Of course, these resonances do not couple in any physical way to the exterior, and resulting observables, but they might influence internal wave dynamics and late-time tails [164], how semi-classical fields diverge or remain finite near the singularity and possible holographic duals or quantum-gravity boundary conditions [165]. Thus, these resonances are physically meaningful in BH interior physics, especially if we consider how fields behave near the singularity or how quantum corrections might modify it.
@@
In quantum gravity, especially in path integral or Euclidean approaches, boundary conditions at singularities or at finite radii are required [175]. Interior resonances could offer a natural selection mechanism for such boundary conditions, e.g., only allowing fields that are smooth or decay in a certain way at $r=0$ [176, 177]. These can even be viewed as “inner quantization conditions” [178], reminiscent of the Bohr-Sommerfeld conditions in semi-classical quantum mechanics [179]. In speculative models where BHs have dual descriptions [180, 181], e.g., fuzzballs [127, 182, 183], firewalls [184], and holographic duals [185, 186, 187], the interior degrees of freedom should encode information about microstates [188]. Interior resonant modes might act like a spectral fingerprint of interior microstructure, even if the actual microstates are unknown. So, even though these modes are not directly observable from outside the BH, they can influence the behavior of semi-classical fields, provide natural regularization or extension criteria, guide the construction of effective theories with modified interiors, or be incorporated into quantum corrections, e.g., via mode sum regularization [189].
@@ Acknowledgements.
We thank Valentin Boyanov, Vitor Cardoso, Nicola Franchini, David Hilditch, Kostas D. Kokkotas, Adrien Kuntz, Jorge Patiño, Hannes Rüter and Sebastian Völkel for helpful discussions. This work was supported by the Israel Science Foundation, grant 531/22. K.D. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., under the Scientific Employment Stimulus – Individual Call – Grant No. 2023.07417.CEECIND/CP2830/CT0008. R.B. acknowledges financial support provided by FCT – Fundação para a Ciência e a Tecnologia, I.P., through the ERC-Portugal program Project “GravNewFields”. K.D. and R.B. also acknowledge financial support provided by the Fundação para a Ciência e Tecnologia (FCT), Portugal, for the financial support to the Center for Astrophysics and Gravitation (CENTRA/IST/ULisboa) through grant No. UID/PRR/00099/2025 and grant No. UID/00099/2025. This project has received funding from the European Union’s Horizon MSCA-2022 research and innovation programme “Einstein Waves” under grant agreement No. 101131233.
@@ B.2 Behavior under the Chandrasekhar-Darboux map
Thus, the possible fate of isospectrality can be decided from the requirement of regularity in the vicinity of $r=0$. If $\psi_{\rm RW}\sim r^{3}$, then
@@
Therefore, although the Regge-Wheeler and Zerilli equations are Darboux-related, their boundary-value problems are may not be isospectral in the BH interior. Hence, axial and polar modes inside Schwarzschild BHs may not share the same spectra.